OpenStudy (anonymous):

laplace transform Obtain an expression for the Laplace transform of 3x"(t )+ 2x'(t )− 2x(t ) where L{x(t )}= Xs, x(0) = −4 and x'(0) = 3

OpenStudy (anonymous):

apply 2 basic formulas of laplace transform\[L{(f')}=sF(s)-f(0)\]\[L(f'')=s^2F(s)-sf'(0)-f(0)\]where\[F(s)=L(f)\]